1. Three capacitors with individual capacitances of 2 μF, 5 μF, and 10 μF respectively are connected in series with a 12V battery. What are the total capacitance and total charge in the network? 2. Same capacitors in number 1 connected in a parallel. If the combined charge in the network is 50 μC, what are the total voltage and total capacitance in the network?


1
Expert's answer
2022-03-07T12:40:28-0500

Series combination

"\frac{1}{C}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}"

"\frac{1}{C}=\frac{1}{2}+\frac{1}{5}+\frac{1}{10}"

"C=\frac{5+2+1}{10}\mu F=\frac{8}{10}\mu F"

"C=0.8\mu F"

Charge

"Q=CV=0.8\times12\mu F=9.6\mu F"(B) same capacitor connected parallel combination

"C=C_1+C_2+C_3"

"C=(2+5+10)\mu F=17\mu F"

"V=\frac{Q}{C}=\frac{50\times10^{-6}}{17\times10^{-6}}=2.95V"